A teacher wanted to see if social anxiety affects performance on an oral assignment. To examine this, students completed social anxiety scales and were classified into “low,” “average,” and “high” levels of social anxiety. The data presented in the table below are the oral presentation grades (in percentages) earned by the 45 students separated by their social anxiety classification. Conduct the most appropriate statistical analysis to determine whether grades on the oral assignment differed based on level of social anxiety. (Hint: see week 1 presentations)

Oral presentation grade (%)

Low

social anxiety

Average

social anxiety

High

social anxiety

72

80

77

68

85

71

73

73

73

75

81

69

97

73

90

73

91

74

95

73

76

88

73

67

89

71

81

97

75

75

88

78

68

89

90

79

93

89

58

89

88

83

79

75

81

State the independent variable in the scenario, its scale of measurement, and its number of levels.

Calculate and interpret the effect size
Answer #1

The independent variable here in this case level of social anxiety. It's scale of measurement is Ordinal because it is rank type( LOW,AVERAGE AND HIGH)

It has three 3 levels: LOW,AVERAGE AND HIGH

Null hypothesis H0:μ(L) = μ(A) = μ(Η)

ALTERNATIVE HYPOTHESIS Ha: Not all means are same

where \mu (L) is Population mean for Low social anxiety

\mu (A)is Population mean for AVERAGE social anxiety

\mu (H) is Population mean for HIGH social anxiety

By using EXCEL WE HAVE

Anova: Single Factor Count Sum Sum SUMMARY Groups Low social anxiety Average social anxiety High social anxiety 15 1265 1195

Test statistic F= MS (BETWEEN)/MS( Within)

F= 340.87/70.36

F= 4.84

P value=0.012788

Since P value smaller than level of significance 0.05 therefore SIGNIFICANT

Decision: REJECT NULL HYPOTHESIS H0.

Conclusion: We have sufficient evidence to conclude that grades on the oral assignment differed based on level of social anxiety at 0.05 level of significance.

Effect size= h2 = SSbetween / SStotal

=681.73/3638.8

=0.1874

0.1874 is the proportion of the total variance that is attributed to an effect ( level of social anxiety)